OpenStudy (anonymous):

Which is a possible solution of the equation y = 2x + 8? (0, 0) (1, 11) (2, 10) (3, 14)

OpenStudy (agent47):

1) Is 0=2*0+8 valid? 2) Is 11 = 2*1+8 valid? etc..

OpenStudy (agent47):

3) Is 10 = 2*2+8 valid? 4) Is 14 = 2*3+8 valid?

OpenStudy (anonymous):

got it :) THANK YOU

OpenStudy (agent47):

np